A LIDAR employs a Q-switched laser providing a 10-ns pulse width and the preamplifier and logconverter circuits have a passband of 10 MHz. What is the distance resolution we can achieve?
Answer #1

Distance resolution =cT/2

C=3*10^8 m/sec

T=10nsec

Distance resolution = (3*10^8 x 10*10^-9)/2

=1.5meters